is there is any easy formula 2 find hyberdization

Grade:11

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Ankit

we have a formula

first we find n=1/2[group no central atom +no of  atoms attached to central metal atom (like H , Cl, F..except O) + magnitude of negative charge - magnitude of positive charge]

if n=2 hybridisation is sp

n=3 hybridisation is sp2

n=4 hybridisation is sp3

n=5 hybridisation is sp3d etc..

for ex in BeCl2 n=1/2[2+2]=1/2(4)=2

so hybridisation is SP like that

some exceptions are there but  FOR MOST OF THE COMPOUNDS WE GET HYBRIDISATION

 

All the best

AKASH GOYAL
AskiitiansExpert-IITD

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.



vikas askiitian expert
13 years ago

use HON rule to find hybridisation of any type of molecule

 according to this rule 2x=number of electrons in valence shell -charge present (if any) +number of monovalent atoms(H,Cl etc)

 if x=2 then hybridisation is sp ,if x is 3 then sp2 , if 3 then sp3 and so on

 example H3O+

  2x = 6(number of valence electrons)- (+1)charge +3(monovalent atoms)

        =6-1+3

      x  =4  (sp3 hybridisation)

 approve my ans if u like